OpenStudy (anonymous):

lim (sqrt(x))(lnx) x->0

OpenStudy (anonymous):

that was wrong rewrite as \[\frac{\ln(x)}{\frac{1}{\sqrt{x}}}\] and use l'hopital

OpenStudy (anonymous):

why is the limit as x->0 of lnx negative infinity?

OpenStudy (anonymous):

do you know what the graph of \(y=\ln(x)\) looks like?

OpenStudy (anonymous):

i guess the answer depends on the definition of \(\ln(x)\) if the definition is the inverse of the exponential function \(e^x\) then it is minus infinity since as x to so \(-\infty\), \(e^x\) goes to zero